Consider the probability distribution for a random variable X.variable X.

XX 33 44 55 66 77
P(X)P(X) 0.150.15 0.100.10 0.200.20 0.250.25 0.300.30

Find the probability P(X≤5.5).probability P(X≤5.5).

Use decimal notation. Give your answer to two decimal places.

P(X≤5.5)=
Answer #1

solution:

using probability distribution properties

given table

XX 33 44 55 66 77
P(X)P(X) 0.150.15 0.100.10 0.200.20 0.250.25 0.300.30

less than equal to so

P(X≤5.5)= p(x=5)+ p(x=4)+ p(x=53) = 0.200+0.100+0.150=0.45

probability=0.45